OpenStudy (anonymous):

k = 4; f(x) = 2x3 - 2x2 - 3x - 5; Lower bound?

OpenStudy (mathmale):

Please explain what your 'k' represents and how it impacts upon f(x). Have you copied down every part of the question at hand?

OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 - 4x3 + 2x2 - 4x + 1; d(x) = x2 + 1

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Is this a new problem, separate from and independent of your original problem (starting with k=4)?

OpenStudy (mathmale):

Regarding the division problem you present here, looks as though you're going to need to do this division using long division. Comfortable with that?